OpenStudy (anonymous):

Solve the system for x: y = 3x + 1 y = 2x² - 5

OpenStudy (campbell_st):

they are both equal to y so equate the equations \[3x + 1 = 2x^2 - 5\] then rewriting the equation you get \[2x^2 -3x - 6 = 0\] the the solution can be found using the general quadratic formula \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] where a = 2, b = -3 and c = -6 hope this helps

OpenStudy (anonymous):

So i got (-3 plusorminus sqrt -23)/4

OpenStudy (campbell_st):

oops slight numerical error it should be \[x = \frac{3 \pm \sqrt{9 + 48}}{4}\]

OpenStudy (anonymous):

Oh. (3 plusorminus sqrt -63)/4

OpenStudy (campbell_st):

it looks like \[x = \frac{-(-3) \pm \sqrt{(-3)^2 - 4 \times 2 \times -6}}{2 \times 2}\] squaring a negative gives a positive... so it becomes \[x = \frac{3 \pm \sqrt{57}}{4}\]

OpenStudy (anonymous):

I'm so off my game today XD

OpenStudy (campbell_st):

thats ok... i find the GQF a bit of a pain... hope this helps.